Show that the points (-2,5), (-2,-1) and (4,-1) all lie on a circle whose center is at (1,2). Find the length of the radius.

Thank You So Much! God Bless!

Answer by stanbon(75887)

Show that the points (-2,5), (-2,-1) and (4,-1) all lie on a circle whose center is at (1,2). Find the length of the radius.
-----------------------------
Find the distance of each of those points from (1,2)
Example::
d = sqrt[(2-5)^2+(1--2)^2] = sqrt[3^2+3^2] = sqrt(18) = 3sqrt(2)
-----
d = sqrt[(2--1)^2+(1--2)^2] = sqrt{3^2 + 3^2] = sqrt(18) = 3sqrt(2)
etc.
Cheers,
Stan H.
-------------